OpenStudy (anonymous):

Solve equation: tan^2x-2secx give possible solutions

OpenStudy (anonymous):

what's the equation?

OpenStudy (anonymous):

\[\tan ^{2}\theta-2\sec \theta=2\]

OpenStudy (anonymous):

\(\large tan^2\theta-2sec\theta=2 \) \(\large tan^2\theta-2sec\theta-2=0 \) \(\large [tan^2\theta-2sec\theta-2=0]\cdot cos^2\theta \) \(\large sin^2\theta-2cos\theta-2cos^2\theta=0 \) \(\large sin^2\theta-2cos\theta-2(1-sin^2\theta)=0 \) \(\large sin^2\theta-2cos\theta-2+2sin^2\theta=0 \) \(\large 3sin^2\theta-2cos\theta-2=0 \) can you take it from here? it's now a trig equation in quadratic form...

OpenStudy (anonymous):

why did you multiply everything but cos^2theta

OpenStudy (anonymous):

I think you'd be better off making the sub.s tan^2 = sec^2-1

OpenStudy (anonymous):

you get sec^2 x -2sec x -3 =0 (sec x -3)(sec x+1) =0

OpenStudy (anonymous):

okay, i set everything to 0?

OpenStudy (anonymous):

sorry... in the third line i should've changed the sin^2 to cos^2: \(\large sin^2\theta-2cos\theta-2cos^2\theta=0 \) \(\large (1-cos^2\theta)-2cos\theta-2cos^2\theta=0 \) \(\large 1-2cos\theta-3cos^2\theta=0 \) \(\large 3cos^2\theta+2cos\theta-1=0 \) that's better....

OpenStudy (anonymous):

*4th line, not 3rd line

OpenStudy (anonymous):

okay that makes more sense now